Suppose the CPI in 1920 was 17 and 240 in 2016

a.) calculate the rate of inflation form 1920 to 2016, to the nearest integer

b.) calculate an equivalent income in 2016 for an income of $36,000 in 1920, to the nearest dollar
Answer #1

Rate of Inflation from 1920 to 2016 is given as below

r=(240/17)^(1/(2016-1920)-1=0.02796

Rate of Inflation =2.796%

Ans for b)

Equivalent Income=36000*(1.02796)^(2016-1920)=$508,181.91=$508,182